# Excited $\Omega$ hyperon in charmful $\Omega_{b}$ weak decays

### v1

@@ IV Discussions and conclusion
The branching fraction of $\Omega_{b}^{-}\to J/\psi\Omega^{-}$, on the order of $10^{-5}-10^{-3}$, indicates inconclusive calculations from Gutsche:2018utw ; Fayyazuddin:2017sxq ; Hsiao:2021mlp ; Cheng:1996cs ; Rui:2023fiz . On the other hand, the constituent quark model has been applied to $\Omega_{c}^{0}$ decays Wang:2022zja , where ${\cal B}(\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-})/{\cal B}(\Omega_{c}^{0}\to\pi^{+}\Omega^{-})=0.21$ agrees with the experimental value of $0.220\pm 0.059\pm 0.035$ Belle:2021gtf , suggesting its applicability to $\Omega_{b}^{-}$ decays.
@@
It should be pointed out that our calculation of $\Omega_{b}^{-}\rightarrow\Omega^{-}J/\psi$ depends on $\alpha$ and $\beta$. To test the sensitivity, we depict $\Gamma(\Omega_{b}^{-}\rightarrow J/\psi\Omega^{-})$ as a function of $\alpha$ and $\beta$, within the parameter space: $0.7~{}{\rm GeV}<\alpha<1.1~{}{\rm GeV}$ and $0.48~{}{\rm GeV}<\beta<0.55~{}{\rm GeV}$. In Fig. 2, it can be seen that $\Gamma(\Omega_{b}^{-}\rightarrow J/\psi\Omega^{-})$ is more sensitive to $\beta$, varying from 0.7 to 1.1 times $10^{-16}$ GeV. However, it remains close to the central value: $0.8\times 10^{-16}$ GeV (see Table 1).
There are two excited $1P$-wave $\Omega$ states in the $\Omega$ hyperon spectroscopy: $\Omega(1^{2}P_{1/2^{-}})^{-}$ and $\Omega(1^{2}P_{3/2^{-}})^{-}$. It is interesting to note that the newly observed $\Omega(2012)^{-}$ state is more likely to be assigned to the latter: $\Omega(1^{2}P_{3/2^{-}})$, based on the fact that the measured mass and decay width are consistent with the quark model predictions Xiao:2018pwe ; Liu:2019wdr ; Aliev:2018yjo ; Aliev:2018syi ; Polyakov:2018mow .
Based on this assignment, we obtain ${\cal B}(\Omega_{b}^{-}\to J/\psi\Omega(2012)^{-})=2.7\times 10^{-4}$. Specifically, ${\cal B}(\Omega_{b}^{-}\to J/\psi\Omega(2012)^{-})/{\cal B}(\Omega_{b}^{-}\to J/\psi\Omega^{-})=1.2$, which clearly avoids the uncertain fragmentation fraction $f_{\Omega_{b}}$, making it beneficial for experimental examination. In fact, $\Omega(2012)^{-}$ has been observed Belle:2018mqs and reconfirmed in $\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-},\Omega(2012)^{-}\to\Xi^{0}K^{-}$ Belle:2021gtf . Likewise, this can be reconfirmed from the resonant $\Omega_{b}^{-}$ decays, such as $\Omega_{b}^{-}\to J/\psi\Omega(2012)^{-}$, followed by the subsequent decay $\Omega(2012)^{-}\rightarrow\Xi^{0}K^{-}$. In particular, $\Omega(2012)\to\Xi K$ has a branching fraction of around $90\%$ Zhong:2022cjx . The $\Omega(1^{2}P_{1/2^{-}})$ is a typically missing $\Omega$ hyperon, with a predicted mass around 1950 MeV Engel:2013ig ; Capstick:1986ter ; Faustov:2015eba ; Liu:2019wdr . In our evaluation, ${\cal B}(\Omega_{b}^{-}\to J/\psi\Omega(1^{2}P_{1/2^{-}})^{-})=1.1\times 10^{-4}$ indicates that $\Omega(1^{2}P_{1/2^{-}})^{-}$ has a smaller but compatible production rate to its $1P$-wave cousin state.
Figure 3: The decay widths of $\Omega_{b}^{-}J/\psi\Omega^{*-}$ as a function of $M_{\Omega^{*}}$, with the range from the theoretical estimations.
@@
In accordance with the masses and wave functions from Ref. Liu:2019wdr , we calculate $\Omega_{b}^{-}\to J/\psi\Omega(1D)^{-}$. Our the results are listed in Table 1, where $\Omega(1^{2}D_{5/2^{+}},1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}},1^{4}D_{7/2^{+}})$ states have production rates to be 0.6, 0.7, 1.0, 0.5, respectively, relative to their ground-state counterpart. We highlight that ${\cal B}(\Omega_{b}^{-}\to J/\psi\Omega(1^{4}D_{5/2^{+}})^{-})=2.1\times 10^{-4}$ is sufficiently large for a promising establishment of $\Omega(1^{4}D_{5/2^{+}})^{-}$ through $\Omega_{b}^{-}\to\Omega(1^{4}D_{5/2^{+}})^{-}J/\psi$, followed by $\Omega(1^{4}D_{5/2^{+}})^{-}\to\Xi(1530)^{0}K^{-}$, where ${\cal B}(\Omega(1^{4}D_{5/2^{+}})\to\Xi(1530)K)$ is as large as $80\%$ Liu:2019wdr ; Xiao:2018pwe .
As for the existing $2S$-wave $\Omega^{*}$ hyperon states, $\Omega(2^{2}S_{1/2^{+}})$ and $\Omega(2^{4}S_{3/2^{+}})$, we obtain ${\cal B}(\Omega_{b}^{-}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}})^{-})=(1.2,2.4)\times 10^{-4}$. Remarkably, in the current $\Omega$ spectroscopy studies, $\Omega(2^{4}S_{3/2^{+}})^{-}$ in $\Omega_{b}^{-}\to J/\psi\Omega^{*}$ is the only one that can have a production rate larger than that of its ground state counterpart.
@@
In conclusion, we have calculated the sextet $b$-baryon decays $\Omega_{b}^{-}\to J/\psi\Omega^{(*)-}$ using the constituent quark model. We found that ${\cal B}(\Omega_{b}^{-}\to J/\psi\Omega^{-})=2.1\times 10^{-4}$, which is consistent with the previous studies. With $\Omega(2012)^{-}$ identified as $\Omega(1^{2}P_{3/2^{-}})$, we found that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012)^{-})=2.7\times 10^{-4}$, which is compatible with ${\cal B}(\Omega_{b}^{-}\to J/\psi\Omega^{-})$. Additionally, the production rates of the $\Omega(1^{2}D_{5/2^{+}},1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}},1^{4}D_{7/2^{+}})$ states have been calculated to be 0.6, 0.7, 1.0, 0.5, respectively, relative to their ground-state counterpart. We also found that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}})^{-})=(1.2,5.1)\times 10^{-4}$, which is promising for measurement by LHCb. Since we have demonstrated that our calculations are insensitive to the parameter inputs and uncertainties arising from the $\Omega^{*}$ masses, this provides a suitable test-bed to investigate the $\Omega$ hyperon spectroscopy.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grants No.12205026, No.12175065 and No.12235018), and Applied Basic Research Program of Shanxi Province, China under Grant No. 202103021223376. YKH was supported by NSFC (Grants No. 12175128 and No. 11675030). JW was supported by General Programs of Changzhi College.

### v4

@@ IV Discussions and conclusion
The branching fraction of $\Omega_{b}\to J/\psi\Omega$, on the order of $10^{-5}-10^{-3}$, indicates inconclusive calculations from Gutsche:2018utw ; Fayyazuddin:2017sxq ; Hsiao:2021mlp ; Cheng:1996cs ; Rui:2023fiz . On the other hand, the constituent quark model has been applied to $\Omega_{c}^{0}$ decays Wang:2022zja , where ${\cal B}(\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-})/{\cal B}(\Omega_{c}^{0}\to\pi^{+}\Omega^{-})=0.21$ agrees with the experimental value of $0.220\pm 0.059\pm 0.035$ Belle:2021gtf , suggesting its applicability to $\Omega_{b}$ decays.
@@
It should be pointed out that our calculation of $\Omega_{b}\rightarrow\Omega J/\psi$ depends on $\alpha$ and $\beta$. To test the sensitivity, we depict $\Gamma(\Omega_{b}\rightarrow J/\psi\Omega)$ as a function of $\alpha$ and $\beta$, within the parameter space: $0.4~{}{\rm GeV}<\alpha<0.5~{}{\rm GeV}$ and $0.48~{}{\rm GeV}<\beta<0.55~{}{\rm GeV}$. In Fig. 2, it can be seen that $\Gamma(\Omega_{b}\rightarrow J/\psi\Omega)$ is more sensitive to $\beta$, varying from 3.0 to 4.2 times $10^{-16}$ GeV. However, it remains close to the central value: $3.6\times 10^{-16}$ GeV (see Table 3).
There are two excited $1P$-wave $\Omega$ states in the $\Omega$ hyperon spectroscopy: $\Omega(1^{2}P_{1/2^{-}})$ and $\Omega(1^{2}P_{3/2^{-}})$. It is interesting to note that the newly observed $\Omega(2012)$ state is more likely to be assigned to the latter: $\Omega(1^{2}P_{3/2^{-}})$, based on the fact that the measured mass and decay width are consistent with the quark model predictions Xiao:2018pwe ; Liu:2019wdr ; Aliev:2018yjo ; Aliev:2018syi ; Polyakov:2018mow .
Based on this assignment, we obtain ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))=1.1\times 10^{-3}$. Specifically, ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))/{\cal B}(\Omega_{b}\to J/\psi\Omega)\simeq 1.3$, which clearly avoids the uncertain fragmentation fraction $f_{\Omega_{b}}$, making it beneficial for experimental examination. In fact, $\Omega(2012)$ has been observed Belle:2018mqs and reconfirmed in $\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-},\Omega(2012)^{-}\to\Xi^{0}K^{-}$ Belle:2021gtf . Likewise, this can be reconfirmed from the resonant $\Omega_{b}$ decays, such as $\Omega_{b}\to J/\psi\Omega(2012)$, followed by the subsequent decay $\Omega(2012)\rightarrow\Xi^{0}K^{-}$. In particular, $\Omega(2012)\to\Xi K$ has a branching fraction of around $90\%$ Zhong:2022cjx . The $\Omega(1^{2}P_{1/2^{-}})$ is a typically missing $\Omega$ hyperon, with a predicted mass around 1950 MeV Engel:2013ig ; Capstick:1986ter ; Faustov:2015eba ; Liu:2019wdr . In our evaluation, ${\cal B}(\Omega_{b}\to J/\psi\Omega(1^{2}P_{1/2^{-}}))=4.2\times 10^{-4}$ indicates that $\Omega(1^{2}P_{1/2^{-}})$ has a smaller but compatible production rate to its $1P$-wave cousin state.
Figure 3: The decay widths of $\Omega_{b}\rightarrow J/\psi\Omega^{*}$ as a function of $M_{\Omega^{*}}$, with the range from the theoretical estimations.
@@
In accordance with the masses from Ref. Liu:2019wdr , we calculate $\Omega_{b}\to J/\psi\Omega(1D)$. Our the results are listed in Table 3, where $\Omega(1^{2}D_{5/2^{+}},1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}},1^{4}D_{7/2^{+}})$ states have production rates to be 0.5, 0.6, 0.8, 0.4, respectively, relative to their ground-state counterpart. We highlight that ${\cal B}(\Omega_{b}\to J/\psi\Omega(1^{4}D_{5/2^{+}}))=7.2\times 10^{-4}$ is sufficiently large for a promising establishment of $\Omega(1^{4}D_{5/2^{+}})$ through $\Omega_{b}\to\Omega(1^{4}D_{5/2^{+}})J/\psi$, followed by $\Omega(1^{4}D_{5/2^{+}})\to\Xi(1530)^{0}K^{-}$, where ${\cal B}(\Omega(1^{4}D_{5/2^{+}})\to\Xi(1530)K)$ is as large as $80\%$ Liu:2019wdr ; Xiao:2018pwe .
As for the existing $2S$-wave $\Omega^{*}$ hyperon states, $\Omega(2^{2}S_{1/2^{+}})$ and $\Omega(2^{4}S_{3/2^{+}})$, we obtain ${\cal B}(\Omega_{b}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}}))=(4.5,20)\times 10^{-4}$. Remarkably, in the current $\Omega$ spectroscopy studies, $\Omega(2^{4}S_{3/2^{+}})$ in $\Omega_{b}\to J/\psi\Omega^{*}$ is the only one that can have a production rate larger than that of its ground state counterpart.
@@
In conclusion, we have calculated the sextet $b$-baryon decays $\Omega_{b}\to J/\psi\Omega^{(*)}$ using the constituent quark model. We found that ${\cal B}(\Omega_{b}\to J/\psi\Omega)=8.8\times 10^{-4}$, which is consistent with the previous studies. With $\Omega(2012)$ identified as $\Omega(1^{2}P_{3/2^{-}})$, we found that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))=1.1\times 10^{-3}$, which is compatible with ${\cal B}(\Omega_{b}\to J/\psi\Omega)$. Additionally, the production rates of the $\Omega(1^{2}D_{5/2^{+}},1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}},1^{4}D_{7/2^{+}})$ states have been calculated to be 0.5, 0.6, 0.8, 0.4, respectively, relative to their ground-state counterpart. We also found that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}}))=(4.5,20)\times 10^{-4}$, which is promising for measurement by LHCb. Since we have demonstrated that our calculations are insensitive to the parameter inputs and uncertainties arising from the $\Omega^{*}$ masses, this provides a suitable test-bed to investigate the $\Omega$ hyperon spectroscopy.
@@ Acknowledgements
We are very grateful to the referee for pointing out the incorrect calculation of the color factor, and to Qiang Zhao and Di Wang for their very helpful discussions regarding the color factor. This work was supported by the National Natural Science Foundation of China (Grants No.12205026, No.12175065 and No.12235018), and Applied Basic Research Program of Shanxi Province, China under Grant No. 202103021223376. YKH was supported by NSFC (Grants No. 12175128 and No. 11675030). JW was supported by General Programs of Changzhi College.